OpenStudy (anonymous):

Among 40 people, 25 like black, 30 like red, 35 like white, at least how many people like all 3 colours? Please show your thought process.

OpenStudy (midhun.madhu1987):

 a+b+c+d+e+f+g = 40 a+b+d+e = 25 b+c+d+f = 30 d+e+f+g = 35 In the figure, the number of people who likes all the color is "d". We need to find d. But i am getting the value of d :( You can make a try...

OpenStudy (haseeb96):

(R) + (B) + (W) - O(R union B union W) 30 +25 +35 -40 90-40 50 are those people who like all the three colours

OpenStudy (midhun.madhu1987):

:)

OpenStudy (anonymous):

No correct answer yet.

OpenStudy (haseeb96):

what are the options there?

OpenStudy (midhun.madhu1987):

Is there anything else in the question..?

OpenStudy (inkyvoyd):

stop putting hte pigeons in pigeonholes

OpenStudy (haseeb96):

is there answer of 70 ?

OpenStudy (haseeb96):

O(R interscetion B) =O(R) + O (B) -O(RUB) = 30 +25 - 40 = 55-40 = 15 O(B intersection W ) = O(B) + O(W) - O (BUW) = 25 +35 -40 = 20 O (W intersection with R ) = O(W) + O(R) - O (WUR) = 35 +30 -40 = 25

OpenStudy (haseeb96):

50 or 70 is the correct answer i think 50 is correct one

OpenStudy (anonymous):

At least 15 people have to like both black and red, 25 at least like both red and white, 20 at least like black and white. At least 10 people have to like the same 3 colors. I use the box method. I drew a box and split it into 4 sections each representing 10 people...

OpenStudy (aum):

 Minimum value for x is 10.